# A voice for minorities: diversity in approval-based committee elections under incomplete or inaccurate information

### v1

@@ Abstract
We study diversity in approval-based committee elections with incomplete or inaccurate information. As standard in the literature on approval-based multi-winner voting, we define diversity according to the maximum coverage problem, which is known to be NP-complete, with a best attainable polynomial time approximation ratio of $1-1/\operatorname{\text{e}}$. In the incomplete information model, voters can vote on only a small portion of the candidates. We suggest a greedy algorithm and a local search algorithm that query voters and use the query responses to approximate the total population’s opinion. For both algorithms, we prove an upper bound on the number of queries required to get a close to $(1-1/\operatorname{\text{e}})$-approximate solution with high probability. We also provide a lower bound for the query complexity of non-adaptive algorithms, that cannot adapt their querying strategy to readily obtained information. In the inaccurate information setting, voters’ responses are corrupted with a probability $p\in(0,\frac{1}{2})$. We provide both an upper and a lower bound for the number of queries required to attain a $(1-1/\operatorname{\text{e}})$-approximate solution with high probability. Finally, using real data from Polis, we see that our algorithms perform remarkably better than the theoretical results suggest, both with incomplete and inaccurate information.
1 Introduction
@@
Consensus has grown among political scientists, democracy practitioners, and decision makers alike, that effective involvement of citizens in policy decisions should be regarded as a high priority for democratic institutions at all levels, from local, to national, to regional [18, 24]. [^1] See also initiatives such as “Conference on the Future of Europe” (2021-22) and “European Democracy Shield” (2024). Many well-established democratic innovations [34], such as citizens’ assemblies and participatory budgeting, contribute to this civic participation goal. Yet, it is on digital democracy platforms in particular where research and development efforts are currently concentrating [6, 13], with the aim of providing citizens with effective tools that could support forms of democratic deliberation at scale [28, 25].
@@
A wealth of this type of digital democracy tools have been developed and deployed around the world in the last decade: LiquidFeedback [3], Consul (recipient of the United Nations Public Service Award), YourPriorities, Decidim, Polis [30], to mention a few. [^2] See www.liquidfeedback.org; www.consulproject.org; www.citizens.is; www.decidim.org; www.pol.is. In different forms, and to a smaller or larger extent, all these applications provide users with the freedom to contribute statements, ideas, arguments to public deliberations on issues relevant to their communities and enable them to express their own opinions on the statements/ideas/arguments expressed by others. Supporting such processes online in large groups necessarily involves the implementation of algorithmic solutions for information-processing problems. One such problem concerns how to feed back to users information that can effectively summarize the current state of the deliberation: how does ‘the group’ think about the issues being deliberated upon?
@@
One approach, which is followed for instance in Polis, to aggregate the current state of the deliberation into a comprehensive summarization of it, is by selecting sets of statements proposed by users, based on how much support the statements have elicited from other users. As [15] already notice, this approach can be conceptualized as an approval-based multi-winner voting problem [11, 21], as studied in computational social choice [2]: users ‘vote’ for other users’ statements by approving or disapproving them. Unlike in standard multi-winner approval voting, however, information about who supports which statements is sparse in online deliberation summarization, as no user can possibly express whether they agree or disagree with every single statement contributed by their peers. For the same reason, voters can become inaccurate in representing their true beliefs, when having to answer many questions. So, such a deliberation summarization problem can be thought of as an approval-based multi-winner voting problem in which ballots are incomplete or inaccurate. From this perspective, [15] focused on the problem of designing selection algorithms that could yield summarizations that meet specific forms of proportional representation proposed in the computational social choice literature [1], while querying users’ opinions as efficiently as possible.
@@
Our work builds on the approach put forth in [15], but leaves proportional representation aside to focus on the property of diversity. Diversity has received much less attention than proportional representation in the multi-winner voting literature [21]. Yet, we argue diversity is especially important for deliberation summarization. Firstly, diverse summarizations would guarantee higher inclusivity, an attribute one could desire in and of itself. Secondly, a diverse summary gives a broad, informative view of ‘the people’s’ opinion. This can be valuable especially in deliberation summarization contexts, where the number of candidates may be very large. Thirdly, cognitive diversity is argued, e.g., by [18], to also be of epistemic advantage from a crowd-wisdom perspective. This is relevant because the selected statements typically serve as a basis for further discussions. We start from the established formalization of diversity in multi-winner voting based on the Chamberlin-Courant [7] score and the Maximum Coverage problem [26, 32]. Our focus thus lies on studying the computational problem of constructing summarizations that maximize the Chamberlin-Courant score, while querying users’ opinions as efficiently as possible.
2 Related work and our contribution
Diversity in multi-winner voting and the Maximum Coverage problem
Diversity in multi-winner voting is commonly defined following the characterization in the seminal work of Chamberlin and Courant in 1983 on representative deliberations and representative decisions [7]. Their most general definition uses voters’ ordinal preferences and, for any given committee, a voter’s representative in the committee is the member who ranks highest in the voter’s preference order. The committee score is obtained by summing over the individual voters’ scores, attributed to them according to the position of their representative in their preference order. In the case of approval ballots, maximizing the Chamberlin-Courant score simplifies to selecting a set of candidates that guarantees for as many voters as possible at least one candidate they approve of. In 2008, [29] have proven, using a reduction from Exact 3-Cover, the Chamberlin-Courant decision problem is NP complete, even for the version with approval ballots. Denoting candidates by the set of their approving voters, the approval-based problem is also equivalent to selecting a set of candidates that maximizes the coverage of the set of voters. This one-to-one correspondence between Chamberlin-Courant score-based diversity in approval-based multi-winner voting, and the Maximum Coverage problem, was first studied in [32].
An instance of the Maximum Coverage problem, Max Cover in short, consists of a set $V$ of $n$ elements, a weight function $w:V\to\mathbb{R}_{>0}$, a family $C\subseteq\mathcal{P}(V)$ of $m$ subsets, and a number $k$. The goal then is to select $k$ sets $c_{i}\in C$ so that $w(\cup_{i=1}^{k}c_{i})$ is maximized. A generalized version of Max Cover, where the aim is to maximize, by electing a set of subsets, any submodular set function [^3] For $C$ a finite set, a set function $f:\mathcal{P}(C)\to\mathbb{R}$ is submodular if, for any $W_{1},W_{2}\subseteq C$, we have $f(W_{1})+f(W_{2})\geq f(W_{1}\cup W_{2})+f(W_{1}\cap W_{2})$., was already studied in 1978, by [26], who proved that a greedy algorithm is $(1-1/\operatorname{\text{e}})$-approximate. However, it was not until twenty years later that the Maximum Coverage problem itself was introduced, by [16], who note NP-completeness of the unit-weight decision problem formulation of Max Cover, by reduction from Set Cover, also describe the greedy algorithm, and prove the approximation ratio of $1-1/\operatorname{\text{e}}$. In the same year, [10] gave yet another proof of the approximation ratio of the greedy algorithm, but more importantly, using a reduction from approximating Max 3SAT-5, Feige proved that for any $\epsilon>0$, Max Cover cannot be approximated in polynomial time within a ratio of $1-1/\operatorname{\text{e}}+\,\epsilon$, unless P=NP; the simple greedy algorithm is thus optimal.
@@ 2 Related work and our contribution
In the literature on multi-winner voting, [32] make boundedness assumptions on the amount of sets an element appears in. The algorithm optimizes over the $\lceil\frac{2pk}{1-\beta}+k\rceil$ largest sets, outputs a $\beta$-approximate committee in time $\text{poly}(n,m)\cdot\binom{2pk/(1-\beta)+k}{k}$, under the assumption of a maximum number $p$ of sets any element may appear in. Under the assumption that each element appears in at least $p$ sets, the standard greedy algorithm attains an improved approximation ratio of $1-\text{e}^{-\max\left(\frac{pk}{m},1\right)}$.
Finally, a slight modification of the problem was studied by [12], who present a combinatorial $(1-1/\operatorname{\text{e}})$-approximation algorithm for Max Cover over a matroid constraint, using a non-oblivious local search that considers improvements of size 1 only. Non-oblivious refers to the algorithm using an auxiliary objective function to evaluate solutions. The paper also shows that, over a matroid constraint, oblivious local search, guided only by the problem’s objective function, achieves an approximation ratio of only $\frac{k-1}{2k-\ell-1}$, when improvements of size $\ell$ are considered. A greedy approach strands at ratio $1/2$.
@@
Single-winner voting with incomplete information has received some attention over the years in the computational social choice literature [5]. More recently, [15] studied proportionality in approval-based committee elections under incomplete information. They present a version of local search PAV that queries voters and, using $\mathcal{O}(mk^{6}\log k\log m)$ queries, finds a solution that satisfies Extended Justified Representation [1] and Optimal Average Satisfaction (based on the notion of average satisfaction, introduced by [33]) with high probability. Their result extends to an $\alpha$-approximate version of both axioms, in which case the query complexity decreases by a factor $k^{3}$. The same study also presents a lower bound of $\Omega(m^{11})$ for the query complexity of non-adaptive algorithms, that can not adapt their querying strategy to readily obtained information. In the field of fair division of indivisible items, [19] study the number of queries required to find a round-robin allocation. When voters’ responses are corrupted with a probability $p\in(0,1/2)$, the query complexity is lower bounded by $\Omega(nm\log(1/\delta)+m\log(m/\delta))$ and upper bounded by $\mathcal{O}(nm\log(m/\delta))$, in both of their two query models.
@@
We study diversity in approval-based committee elections when information elicited from voters is incomplete or inaccurate. We measure diversity of a solution with its unit weight Max Cover, or equivalently, approval-based Chamberlin-Courant, score. For a large part, our work combines the two lines of research outlined above, and is based in particular on [12], [15] and [19]. We make four main contributions. First, we adapt the local search algorithm for Max Cover over a matroid constraint studied by [12] to an incomplete information setting, and prove an upper bound on the number of queries required to attain a $(1-1/\operatorname{\text{e}})\gamma$-approximate solution with high probability, where $\gamma$ is a parameter in $(0,1)$ (Theorem 4.10). For the original Max Cover problem (i.e. not over a matroid constraint) we study a greedy algorithm and also prove an upper bound for the number of queries required for execution (Theorem 4.4). Second, we prove, for any $\epsilon>0$ and $\gamma\in(\frac{1}{2},1)$, a lower bound for the query complexity of any non-adaptive algorithm that is $(1-1/\operatorname{\text{e}})\gamma$-approximate with probability more than $\epsilon$ (Theorem B.1). Our proof follows the proof strategy of [15, Theorem 3.1]. Third, for a setting with inaccurate information, we adapt the proof of Theorem 9 of [19], and establish a lower bound of $\Omega(nm\log(1/\delta))$ for the query complexity of finding a $(1-1/\operatorname{\text{e}})$-approximate solution with probability $1-\delta$ (Theorem 5.2). We attain an upper bound of $\mathcal{O}(nm\log(nm/\delta)$ (Theorem 5.1). Fourth, using real data from Polis, we provide a number of observations suggesting that our algorithms perform considerably better than what our worst-case analysis suggests. The observation appears to hold even in situations in which votes are both incomplete and inaccurate. For an overview of the theoretical results, see Figure 1.
Figure 1: Overview of the theoretical results. For specifics on e.g. range of parameters, or exceptions, see the theorem statements.
3 Preliminaries
@@
An instance of an approval-based committee election problem consists of a set $V$ of $n$ voters, a set $C$ of $m$ candidates, and a natural number $k$. The goal is to elect a committee $W\subseteq C$ of size $k\leq m$, based on the opinions of the voters in $V$. In the context of deliberation summarization, we often even have $k\ll m$. An instance also contains approval information: every voter $i\in V$ approves of a subset of the candidates. Usually, this approval information is known upfront and expressed in the form of an approval set $A_{i}$ for each voter $i\in V$. The sequence of sets $A=(A_{1},\ldots,A_{n})$ is called the approval profile, and we call this setting the perfect information setting. This paper studies incomplete and inaccurate information settings; the respective models are detailed in the corresponding section.
To measure diversity of an elected committee, we use the approval-based Chamberlin-Courant score, which is equivalent to the unit weight Max Cover score:
Definition 3.1 ([7]).
@@
On any approval-based committee election instance $(V,C,A,k)$, the approval-based Chamberlin-Courant (CC) score of a committee $W\subseteq C,\,|W|=k$ is
$\text{CC}(W)=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}_{\{A_{i}\cap W\neq\emptyset\}}(i).$ (1)
@@
In the approval-based Chamberlin-Courant decision problem, the input is an instance $(V,C,A,k)$ and an integer $x$, and the question is whether it is possible to achieve score $\frac{x}{n}$ using $k$ candidates. In the approval-based Chamberlin-Courant optimization problem, the input is an instance $(V,C,A,k)$ and the task is to maximize the score using $k$ candidates.
[…]
Since we study only approval-based committee elections, and the approval-based Chamberlin-Courant problem is equivalent to the unit-weight Maximum Coverage problem, we use both terms interchangeably. Similarly, we refer interchangeably to voters/elements, to candidates/sets, and committees/covers. Whenever we speak about the Chamberlin-Courant problem, the approval-based version is meant, and whenever the Maximum Coverage problem is mentioned, the unit-weight version is meant.
4 Incomplete information
In the incomplete information setting, voters are only able to express their votes (approve/disapprove) on a small portion of the candidates, so we don’t have access to all approval information beforehand. To model this, we assume there exists an underlying true distribution $d$ of approval preferences, and we now write $A:V\to\mathcal{P}(C)$ for a random variable, distributed according to $d$, that maps a voter $i$ to the set $A(i)$ of candidates that voter $i$ approves of. To obtain approval information, an algorithm must draw samples from $d$. We assume that voters are able to respond to at most $t$ approve/disapprove questions at one go, so drawing voter $i$, we obtain, for some set $Q\subseteq C$ for which $|Q|=t$, the set $A(i)\cap Q$. We call the voter-candidate pair $(i,j)$ a query and $Q$ a query set. This means query set $Q$ allows us to perform $t$ queries per queried voter. We are interested in finding a solution to Max Cover in this incomplete information setting. When voters are able to express their vote on all candidates, $nm$ query responses suffice to gather all information, and we will surely find a $(1-1/\operatorname{\text{e}})$-approximate solution, e.g. with a greedy approach. When voters are not able to express their vote on all candidates, the algorithm’s performance might drop. In this section, we present theoretical results on the relation between the approximation ratio and the number of queries required to obtain that ratio. We study two settings. Firstly, the original problem, where we may elect any $k$-sized committee we want, and, secondly, the problem over a matroid constraint, where the elected committee has to be part of a matroid that is defined with the problem instance. For the proof of a lower bound for the query complexity of non-adaptive algorithms in the original problem, we refer to the appendix.
4.1 Upper bound for the unconstrained setting
@@
We prepare the ground by considering the perfect information setting. Algorithm 1 is the greedy algorithm for this setting. We write $\Delta(W,c^{\prime})$ for the increase in Chamberlin-Courant score obtained by adding $c^{\prime}$ to committee $W$. Here, but also in future algorithms, we abuse notation by using $\in$ for lists.
@@
Let $W=[\ ]$ be an empty list;
@@
    For all $c^{\prime}\notin W$, determine $\Delta(W,c^{\prime})$.
@@
    Append $W$ with $\operatorname*{argmax}_{c^{\prime}\notin W}\,\Delta(W,c^{\prime})$.

### v2

# Diverse Committees with Incomplete or Inaccurate Approval Ballots

@@ Abstract
We study diversity in approval-based committee elections with incomplete or inaccurate information. We define diversity according to the Maximum Coverage problem, which is known to be np-complete, with a best attainable polynomial time approximation ratio of $1-1/\operatorname{\text{e}}$. In the incomplete information setting, voters vote only on a small portion of the candidates, and we prove that getting arbitrarily close to the optimal approximation ratio w.h.p. requires $\Omega(m^{2})$ non-adaptive queries, where $m$ is the number of candidates. This motivates studying adaptive querying algorithms, that can adapt their querying strategy to information obtained from previous query outcomes. In that setting, we lower this bound to only $\Omega(m)$ queries. We propose a greedy algorithm to match this lower bound up to log-factors. We prove the same $\tilde{\Theta}(m)$ bound for the generalized problem of Max Cover over a matroid constraint, using a local search algorithm. Specifying a matroid of valid committees lets us implement extra structural requirements on the committee, like quota. In the inaccurate information setting, voters’ responses are corrupted with a small probability. We prove $\tilde{\Theta}(nm)$ queries are required to attain a $(1-1/\operatorname{\text{e}})$-approximation with high probability, where $n$ is the number of voters. While the proven bounds show that all our algorithms are viable asymptotically, they also show that some of them would still require large numbers of queries in instances of practical relevance. Using real data from Polis as well as synthetic data, we observe that our algorithms perform well also on smaller instances, both with incomplete and inaccurate information.
Introduction
@@
Consensus has grown among political scientists, democracy practitioners, and decision makers alike, that effective involvement of citizens in policy decisions should be regarded as a high priority for democratic institutions at all levels, from local, to national, to regional [18, 23]. Currently, it is on digital democracy platforms especially that research and development efforts are concentrating [5, 12], in order to provide citizens with effective civic participation tools [27, 24].
@@
A wealth of these types of digital democracy tools have been developed and deployed around the world in the last decade: LiquidFeedback [3], Consul, YourPriorities, Decidim, Polis [29], to mention a few. [^1] See www.liquidfeedback.org; www.consulproject.org; www.citizens.is; www.decidim.org; www.pol.is In different forms, all these applications allow users to provide free-text input to public deliberations and enable them to express their own opinions on the input of others. Supporting such processes involves the use of algorithms for information-processing problems. One such problem concerns how to effectively summarize the current state of the deliberation: how does ‘the group’ currently think about the issues being deliberated upon?
@@
One approach to this problem, which is followed for instance in Polis, consists of selecting sets of statements proposed by users, based on how much support the statements have elicited from other users. As [14] already noticed, this approach can be conceptualized as an approval-based multi-winner voting problem [10, 21], as studied in computational social choice [2]: users ‘vote’ for other users’ statements by approving or disapproving them. Unlike in standard multi-winner approval voting, however, information about who supports which statements is now sparse, as no user can possibly express whether they agree or disagree with every single statement contributed by their peers. For the same reason, voters can become inaccurate in representing their true beliefs, when having to answer many questions. Such a deliberation summarization problem can thus be thought of as an approval-based multi-winner voting problem in which ballots are incomplete or inaccurate, or both. From this perspective, [14] focused on the problem of designing selection algorithms yielding summarizations that meet specific forms of proportional representation proposed in the computational social choice literature [1], while querying users’ opinions as efficiently as possible.
@@
Our work builds on the approach put forth in [14], but focuses on the property of diversity instead. Diversity has received less attention than proportional representation in the multi-winner voting literature [21]. Yet, we argue that diversity is especially important for deliberation summarization. Firstly, diverse summarizations can guarantee higher inclusivity, an attribute one could desire in and of itself [24]. Secondly, a diverse summary gives a broad, informative view of ‘the group’s’ opinion. This can be valuable especially in deliberation summarization contexts, where the number of candidates may be very large. Thirdly, cognitive diversity is argued, e.g., by [18], to also be of epistemic value from a crowd-wisdom perspective. This is relevant because the selected statements typically serve as a basis for further discussions. We use the established formalization of diversity in multi-winner voting based on the Chamberlin-Courant score [6] and the Maximum Coverage problem [25, 31]. Our focus thus lies on the computational problem of constructing summarizations that maximize the Chamberlin-Courant score, while querying users’ opinions as efficiently as possible.
Related work and contribution
Diversity in multi-winner voting and Max Cover
In multi-winner voting, the diversity of a committee is commonly quantified by the Chamberlin-Courant score [6]. In approval-based multi-winner voting, a voter’s representative in a committee $W$ then is any elected candidate they approve of, and the approval-based Chamberlin-Courant (going forward: Chamberlin-Courant) score of $W$ equals the number of voters that have a representative in $W$. [28] proved that the winner determination problem based on the Chamberlin-Courant score is np-complete. Portraying candidates by their set of approving voters, the problem amounts to selecting a set of candidates that maximizes the ‘coverage’ of the set of voters. This one-to-one correspondence between Chamberlin-Courant score-based diversity in approval-based multi-winner voting, and the unit-weight Maximum Coverage problem in combinatorial optimization, was first studied in [31].
@@
Finally, a generalization of the problem was studied by [11], who present a local search algorithm that is $(1-1/\operatorname{\text{e}})$-approximate for Max Cover over a matroid constraint, where the greedy algorithm strands at ratio $1/2$ in this generalized setting.
@@
Single-winner voting with incomplete information has received some attention over the years in the computational social choice literature, see [4]. More recently, [14] studied proportionality in approval-based committee elections under incomplete information. The paper presents a version of local search PAV that queries voters and, using $O(mk^{6}\log k\log m)$ queries, finds a solution that satisfies Extended Justified Representation [1] and Optimal Average Satisfaction (based on the notion of average satisfaction, introduced by [32]) with high probability. The result extends to an $\alpha$-approximate version of both axioms, in which case the query complexity decreases by a factor $k^{3}$. The same study also presents a lower bound of $\Omega(m^{11})$ for the query complexity of non-adaptive algorithms, that can not adapt their querying strategy information obtained from previous queries. In the field of fair division of indivisible items, [19] study the number of queries required to find a round-robin allocation. When voters’ responses are corrupted with a small probability, the query complexity is in $\Omega(nm+m\log m)$ and $O(nm\log m)$.
@@
We study diversity in approval-based committee elections when information elicited from voters is incomplete or inaccurate. We measure diversity of a solution with its Max Cover, or equivalently, Chamberlin-Courant, score. For a large part, our work combines the two lines of research outlined above, and is based in particular on [11], [14] and [19]. We make three main contributions. First, in a setting with incomplete information, we prove that getting arbitrarily close to the optimal approximation ratio w.h.p. requires $\Omega(m^{2})$ non-adaptive queries (Appendix C). This motivates studying adaptive querying algorithms, that can adapt their querying strategy to information obtained from previous query outcomes. In that setting, we lower this bound to only $\Omega(m)$ queries. We adapt the greedy algorithm to match this lower bound up to log-factors (Theorem 2). We prove the same $\tilde{\Theta}(m)$ bound for the generalized problem of Max Cover over a matroid constraint (Theorem 6). Specifying a matroid of valid committees lets us implement external diversity requirements, like quota. Second, in the inaccurate information setting, we prove that recovering the optimal approximation ratio w.h.p. requires $\tilde{\Theta}(nm)$ queries (Theorem 8). Despite these positive asymptotic bounds, summarized in Table 1, our results do involve sizeable constant overheads for some of our algorithms, which make viability on real-world instance questionable. So, third, using real data from Polis and synthetic data that we produced by adapting methods from [33], we empirically show that our algorithms perform considerably better than our worst-case analysis suggests. Importantly, this appears to hold even in situations in which votes are both incomplete and inaccurate.
Incomplete Inaccurate
Adaptive Non-adaptive $\tilde{\Theta}(nm)$
Matroid No matroid $\Omega(m^{2})$
$\tilde{\Theta}(m)$ $\tilde{\Theta}(m)$
Table 1: Overview of query complexity bounds in $m$ and $n$.
Preliminaries
@@
An instance of an approval-based committee election problem consists of a set $V$ of $n$ voters, a set $C$ of $m$ candidates, and a natural number $k$. The goal is to elect a committee $W\subseteq C$ of size $k\leq m$, based on the opinions of the voters in $V$. We assume, based on observations on online deliberation summarization platforms, that $m\in\Theta(n)$ and $k\ll m$. An instance also contains approval information: every voter $v_{i}\in V$ approves of a subset of the candidates. Usually, this approval information is known upfront and expressed in the form of an approval set $A(i)$ for each voter $v_{i}\in V$. The sequence of sets $A=(A(1),\ldots,A(n))$ is called the approval profile, and we call this setting the perfect information setting. This paper studies incomplete and inaccurate information settings; the respective models are detailed in the corresponding section. To measure diversity of an elected committee, we use the Chamberlin-Courant score.
Definition 1 (Chamberlin and Courant 1983).
@@
On any approval-based committee election instance $(V,C,A,k)$, the Chamberlin-Courant (CC) score of a committee $W\subseteq C,\,|W|=k$ is
$\text{CC}(W)=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}_{\{A(i)\cap W\neq\emptyset\}}(i).$ (1)
@@
In the Chamberlin-Courant decision problem, the input is an instance $(V,C,A,k)$ and an integer $x$, and the question is whether it is possible to achieve score $\frac{x}{n}$ using $k$ candidates. In the Chamberlin-Courant optimization problem, the input is an instance $(V,C,A,k)$ and the task is to maximize the score using $k$ candidates.
We will use the terms Max Cover and Chamberlin-Courant interchangeably. We do the same for voters/elements, to candidates/sets, and committees/covers.
Incomplete information
In the incomplete information setting, the complete approval profile $A=(A(1),\dots,A(n))$ is replaced by a function $A(i,j):=\mathds{1}_{\{c_{j}\in A(i)\}}(i,j)$ which equals 1 in case voter $v_{i}$ approves candidate $c_{j}$ and 0 otherwise. This is a generalization of the original model, as querying all voters on all candidates recovers the complete approval profile. As such, by making $nm$ queries to $A$ we can find a $(1-1/\operatorname{\text{e}})$-approximate solution, e.g. with a greedy approach. In Appendix C, we show that any non-adaptive querying algorithm must make $\Omega(m^{2})$ queries to $A$ to get arbitrarily close to this optimal approximation ratio w.h.p., where non-adaptive means that the sequence of queries made must be specified beforehand. In our setting, $m\in\Theta(n)$, which means $\Omega(m^{2})=\Omega(nm)$, making no improvement over the complete information setting. We can do better studying adaptive querying algorithms, that can adapt their querying strategy to information obtained from previous query outcomes. For such adaptive algorithms we will show that obtaining an approximation ratio arbitrarily close to this optimal ratio w.h.p. requires only $\tilde{\Theta}(m)$ queries to $A$.
To see that we need at least $\Omega(m)$ queries, imagine an instance where only one candidate has any approvals. Any algorithm must select this candidate to attain a positive approximation ratio. In the worst-case, finding this voter w.h.p. requires sampling voters for each of the $m$ candidates. In the remainder of the section we study algorithms that witness this lower bound $\Omega(m)$. We first adapt a greedy algorithm for Max Cover and then do the same for a local search algorithm for the generalized problem of optimizing over a matroid constraint. Both algorithms achieve an approximation ratio arbitrarily close to the optimal ratio w.h.p. using $\tilde{O}(m)$ queries, matching the lower bound (up to log-factors).
Upper bound for the unconstrained setting
@@
We prepare the ground by considering the greedy algorithm for the perfect information setting, Algorithm 1, which achieves the optimal approximation ratio of $1-1/\operatorname{\text{e}}$ [15]. Algorithm 1 elects, in each iteration, the candidate that yields the largest immediate increase in Max Cover score. We write $\Delta(W,c)$ for the increase in Max Cover score obtained by adding candidate $c$ to committee $W$. We give a novel proof of its approximation ratio in Appendix A.
@@
Let $W=\{\}$ be an empty set;
for $i=1,\ldots,k$, do
   Add $c^{\prime}\in\operatorname*{argmax}_{c\notin W}\,\Delta(W,c)$ to $W$.
   end for
ALGORITHM 1 greedy
The greedy query algorithm
We adapt Algorithm 1 to the incomplete information setting. Our strategy is to repeatedly sample sets of voters (of size $\ell\leq n$) uniformly at random, in order to estimate $\Delta(W,c)$ for all $c\notin W$. The algorithm then behaves like the standard greedy algorithm: iteratively selecting the candidate $c$ maximizing this estimate of $\Delta(W,c)$. By carefully bounding the deviation from the true value $\Delta(W,c)$, we can guarantee that our algorithm attains a ratio arbitrarily close to the optimal approximation ratio.
To write this more formally, note that if we ask all voters for their votes on some query set $Q\subseteq C$ (of size $t$), the responses allow us to, for any set $S\subseteq Q$, compute
$p_{S}:=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}_{\{A(i)\cap S\neq\emptyset\}}(i).$ (2)
The value $p_{S}$ equals the Chamberlin-Courant score of $S$, and in this case $\Delta(W,c)=p_{W\cup\{c\}}-p_{W}$, so we can compute $\Delta(W,c)$ with complete information on a query set containing $W\cup\{c\}$. Sampling $\ell$ voters uniformly at random and querying them the set $Q\subseteq C$, we can compute, for every subset $S\subseteq Q$,
$\hat{p}_{S}:=\frac{1}{\ell}\sum_{i=1}^{\ell}\mathds{1}_{\{A(i)\cap S\neq\emptyset\}}(i),$ (3)
that estimates $p_{S}$. We write $\hat{\Delta}(W,c):=\hat{p}_{W\cup\{c\}}-\hat{p}_{W}$ for an estimate of $\Delta(W,c)$ based on $\hat{p}$. Using this approach, we obtain Algorithm 2.
We assume that $\sum_{v_{i}\in V}\mathds{1}_{\{|A(i)|\geq 1\}}(i)\geq k$; there exist at least $k$ voters with a non-empty approval set. We call instances for which this is true, non-trivial; a trivial instance would attain maximal score simply by electing a candidate for each voter $v_{i}$ for which $A(i)\cap C\neq\emptyset$. Theorem 2 gives an upper bound on the number of queries required to run Algorithm 2 on non-trivial instances.
Input: Numbers $n,m,k\in\operatorname{\mathbb{N}}$ with $k\leq m$, set $V$ of $n$ voters, set $C$ of $m$ candidates, query size $t$ with $m\geq t>k$, and $\gamma\in(0,1),\ \delta>0$.
Output: Committee $W\subseteq C$ of size $k$.
Let $W=\{\}$ be an empty set;
@@
    Construct the smallest set of query sets $\mathcal{Q}=\{Q_{i}\}_{i}$ with $Q_{i}\subseteq C$ and $|Q_{i}|=t$ for all $i$, and such that $W\subseteq\bigcap\mathcal{Q}$ and $\bigcup\mathcal{Q}=C$, and present each query set to $\ell$ voters sampled uniformly at random.
@@
    For all $c\notin W$, determine $\hat{\Delta}(W,c)$ as an estimate of $\Delta(W,c)$ using $\ell$ voters and query set $Q$ containing $W\cup c$.
  3. 3.
    Add $c^{\prime}\in\operatorname*{argmax}_{c\notin W}\,\hat{\Delta}(W,c)$ to $W$.